A school boy standing 100m away from a wall, claps his hands repeatedly at regular intervals to coincide exactly with the echoes . If he made 50 claps in the time of 30s. Calculate the speed of sound in air 


1
Expert's answer
2021-04-26T17:09:03-0400
"v=\frac{2L}{T}\\v=\frac{2(100)(50)}{30}=333\frac{m}{s}"